the population of a midwest town decreased from 214,000 to 211,000 in just 2 years. assuming this trend continues, what will population be after an additional 3 years?
thank you for your time and help

Answer by Boreal(15235)

P=Po{1+r}^t, with t in years
211000=214000(1+r)^2
divide by 214000, both sides
0.9860=(1+r)^2
ln of both sides
-0.0141=2 (ln (1+r)
-0.00706=ln (1+r)
raise both to e
0.9930=1+r
r=-0.007
Now P=Po(1-0.007)^5, because it will be 5 years.
=206,614
With these problems, round at the end. I wrote in the number to a few decimal places, but the calculations were done with the original unrounded answer.